# A Synthesis of Green Architectural Tactics for ML-Enabled Systems

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 46th International Conference on Software Engineering - Software Engineering in Society; April 2024; Lisbon, Portugalprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Designing softwareccs: Software and its engineering Software architecturesccs: Social and professional topics Sustainabilityccs: Computing methodologies Machine learning
@@ 4.1. Data-Centric
T1: Apply sampling techniques
Use a smaller subset of the original input dataset
Energy efficiency
(Verdecchia et al., 2022)(Wang et al., 2019)
T2: Remove redundant data
Detect and remove redundant data from the original input data
Energy efficiency
(Ang et al., 2022b)(Dhabe et al., 2021)
T3: Reduce number of data features
Reduce the number of input data features used
Energy efficiency
(Verdecchia et al., 2022)
T4: Use input quantization
Convert input data to smaller precision
Accuracy*
(Abreu et al., 2022)(Kim et al., 2021)
T5: Use data projection
Project data into a lower-dimensional embedding
Performance*
(Rouhani et al., 2016)
@@ 4.2. Algorithm Design
T6: Choose an energy-efficient algorithm
Choose the most energy-efficient algorithm that achieves sufficient level of accuracy
Energy efficiency
(Kaack et al., 2022)
T7: Choose a lightweight algorithm alternative
If possible, choose lighter alternatives of existing algorithms
Energy efficiency
(Sorbaro et al., 2020)
T8: Decrease model complexity
Decrease the complexity of an ML model
Energy efficiency
(Abreu et al., 2020)(Morotti et al., 2021)
T9: Consider reinforcement learning for energy efficiency
Use reinforcement learning to optimize energy efficiency at run time
Energy efficiency
(Kim and Wu, 2020)(Mohammed et al., 2020)
T10: Use dynamic parameter adaptation
Design parameters that are dynamically adapted based on the input data
Energy efficiency
(García-Martín et al., 2021)
T11: Use built-in library functions
Use built-in libraries for ML models if possible
Performance*
(Shanbhag et al., 2022)
@@
T6: Choose an energy-efficient algorithm. Different ML algorithms have different levels of energy consumption and computational power. For example, K-nearest neighbor (KNN) algorithms have a much higher energy consumption than Random Forest (RF) (Verdecchia et al., 2022). High energy consumption does not necessarily mean that the algorithms perform better or achieve higher accuracy levels than low-energy algorithms. Thus, choosing suitable, energy-efficient algorithms that achieve wanted outcomes can reduce the energy consumption of ML models (Kaack et al., 2022).
@@ 4.3. Model Optimization
T12: Set energy consumption as a model constraint
Consider energy consumption as one predetermined parameter for optimizing the ML model
Energy efficiency
(Wang et al., 2021)(Yang et al., 2019)
T13: Consider graph substitution
Replace energy-intensive model parts with similar, but less energy-consuming parts
Energy efficiency
(Wang et al., 2020)
T14: Enhance model sparsity
Reduce the number of model parameters or set their values to zero
Energy efficiency
(Yang et al., 2020b)
T15: Consider energy-aware pruning
Prune neural networks starting from the most energy-intensive layer
Energy efficiency
(Yang et al., 2017)
T16: Consider transfer learning
Use pre-trained ML models for other similar tasks
Energy efficiency
(Jayakodi et al., 2020)(Shanbhag et al., 2022)
T17: Consider knowledge distillation
Use knowledge from a large ML model to train a smaller model
Performance*
(Shanbhag et al., 2022)(Yang et al., 2019)
@@ 4.4. Model Training
T18: Use quantization-aware training
Convert high-precision data types to lower precision during training
Accuracy*
 (Sorbaro et al., 2020; Kim et al., 2021)
T19: Use checkpoints during training
Use checkpoints to avoid a knowledge loss in case of a premature termination
Recoverability*
 (Shanbhag et al., 2022)
T20: Design for memory constraints
Consider possible memory constraints during training
Recoverability*
 (Shanbhag et al., 2022)
@@ 4.5. Deployment
T21: Consider federated learning
Train the model and store data in decentralized devices
Energy efficiency
 (Kim et al., 2021)
T22 Use computation partitioning
Divide computations between a client and a cloud server
Energy efficiency
 (Manasi et al., 2020)
T23: Apply cloud fog network architecture
Use an architecture in which the models are processed between end devices and cloud
Energy efficiency
 (Yosuf et al., 2021)
T24: Use energy-efficient hardware
Use energy-efficient, ML-suitable hardware
Energy efficiency
 (Kaack et al., 2022)
T25: Use power capping
Set energy consumption limits for hardware
Energy efficiency
 (Krzywaniak et al., 2022)
T26: Use energy-aware scheduling
Dynamically optimize the scheduling of ML tasks
Resource utilization*
 (Sun et al., 2020)
T27: Minimize referencing to data
Avoid unnecessary read and write data operations
Energy efficiency
 (Shanbhag et al., 2022)
@@ 4.6. Management
T28: Use informed adaptation
Adapt the model based on informed concept shift
Energy efficiency
(Poenaru-Olaru et al., 2023)
T29: Retrain the model if needed
In case of concept shift, retrain the existing ML model instead of building a new one
Accuracy*
(Poenaru-Olaru et al., 2023)
T30: Monitor computing power
Monitor computing power of an ML model in the long-term
Energy efficiency
(Cao et al., 2021)(Kumar et al., 2020)
@@ 4.7. Targeted Quality Attributes
Energy efficiency
The ability to accomplish a task while minimizing energy consumption
21
Performance
The efficiency with which a task is achieved (e.g., speed, stability)
3
Accuracy
The level of how accurately the algorithm performs specified tasks.
3
Recoverability
The ability to restore and resume normal operations after a failure
2
Resource utilization
The ability to use and allocate resources efficiently
1
@@ Acknowledgements.
We kindly thank our three focus group experts for their valuable time and feedback! Additionally, we thank Iffat Fatima (VU Amsterdam) for her support with integrating the tactics into the AADT. Lewis and Ozkaya’s work was supported by the Department of Defense under Contract No. FA8702-15-D-0002 with Carnegie Mellon University for the operation of the Software Engineering Institute, a federally funded research and development center (DM23-2363).

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Software Engineering in Society; April 14–20, 2024; Lisbon, Portugalbooktitle: Software Engineering in Society (ICSE-SEIS’24), April 14–20, 2024, Lisbon, Portugaldoi: 10.1145/3639475.3640111isbn: 979-8-4007-0499-4/24/04ccs: Software and its engineering Designing softwareccs: Software and its engineering Software architecturesccs: Social and professional topics Sustainabilityccs: Computing methodologies Machine learning
@@ 4.1. Data-Centric
T1: Apply sampling techniques Use a smaller subset of the original input dataset Energy efficiency (Verdecchia et al., 2022)(Wang et al., 2019)
T2: Remove redundant data Detect and remove redundant data from the original input data Energy efficiency (Ang et al., 2022b)(Dhabe et al., 2021)
T3: Reduce number of data features Reduce the number of input data features used Energy efficiency (Verdecchia et al., 2022)
T4: Use input quantization Convert input data to smaller precision Accuracy* (Abreu et al., 2022)(Kim et al., 2021)
T5: Use data projection Project data into a lower-dimensional embedding Performance* (Rouhani et al., 2016)
@@ 4.2. Algorithm Design
T6: Choose an energy-efficient algorithm Choose the most energy-efficient algorithm that achieves sufficient level of accuracy Energy efficiency (Kaack et al., 2022)
T7: Choose a lightweight algorithm alternative If possible, choose lighter alternatives of existing algorithms Energy efficiency (Sorbaro et al., 2020)
T8: Decrease model complexity Decrease the complexity of an ML model Energy efficiency (Abreu et al., 2020)(Morotti et al., 2021)
T9: Consider reinforcement learning for energy efficiency Use reinforcement learning to optimize energy efficiency at run time Energy efficiency (Kim and Wu, 2020)(Mohammed et al., 2020)
T10: Use dynamic parameter adaptation Design parameters that are dynamically adapted based on the input data Energy efficiency (García-Martín et al., 2021)
T11: Use built-in library functions Use built-in libraries for ML models if possible Performance* (Shanbhag et al., 2022)
@@
T6: Choose an energy-efficient algorithm. Different ML algorithms have different levels of energy consumption and computational power. For example, the K-nearest neighbor (KNN) algorithm has much lower energy consumption than the ensemble method Random Forest (RF) (Verdecchia et al., 2022). High energy consumption does not necessarily mean that the algorithms perform better or achieve higher accuracy levels than low-energy algorithms. Thus, choosing suitable, energy-efficient algorithms that achieve wanted outcomes can reduce the energy consumption of ML models (Kaack et al., 2022).
@@ 4.3. Model Optimization
T12: Set energy consumption as a model constraint Consider energy consumption as one predetermined parameter for optimizing the ML model Energy efficiency (Wang et al., 2021)(Yang et al., 2019)
T13: Consider graph substitution Replace energy-intensive model parts with similar, but less energy-consuming parts Energy efficiency (Wang et al., 2020)
T14: Enhance model sparsity Reduce the number of model parameters or set their values to zero Energy efficiency (Yang et al., 2020b)
T15: Consider energy-aware pruning Prune neural networks starting from the most energy-intensive layer Energy efficiency (Yang et al., 2017)
T16: Consider transfer learning Use pre-trained ML models for other similar tasks Energy efficiency (Jayakodi et al., 2020)(Shanbhag et al., 2022)
T17: Consider knowledge distillation Use knowledge from a large ML model to train a smaller model Performance* (Shanbhag et al., 2022)(Yang et al., 2019)
@@ 4.4. Model Training
T18: Use quantization-aware training Convert high-precision data types to lower precision during training Accuracy*  (Sorbaro et al., 2020; Kim et al., 2021)
T19: Use checkpoints during training Use checkpoints to avoid a knowledge loss in case of a premature termination Recoverability*  (Shanbhag et al., 2022)
T20: Design for memory constraints Consider possible memory constraints during training Recoverability*  (Shanbhag et al., 2022)
@@ 4.5. Deployment
T21: Consider federated learning Train the model and store data in decentralized devices Energy efficiency  (Kim et al., 2021)
T22 Use computation partitioning Divide computations between a client and a cloud server Energy efficiency  (Manasi et al., 2020)
T23: Apply cloud fog network architecture Use an architecture in which the models are processed between end devices and cloud Energy efficiency  (Yosuf et al., 2021)
T24: Use energy-efficient hardware Use energy-efficient, ML-suitable hardware Energy efficiency  (Kaack et al., 2022)
T25: Use power capping Set energy consumption limits for hardware Energy efficiency  (Krzywaniak et al., 2022)
T26: Use energy-aware scheduling Dynamically optimize the scheduling of ML tasks Resource utilization*  (Sun et al., 2020)
T27: Minimize referencing to data Avoid unnecessary read and write data operations Energy efficiency  (Shanbhag et al., 2022)
@@ 4.6. Management
T28: Use informed adaptation Adapt the model based on informed concept shift Energy efficiency (Poenaru-Olaru et al., 2023)
T29: Retrain the model if needed In case of concept shift, retrain the existing ML model instead of building a new one Accuracy* (Poenaru-Olaru et al., 2023)
T30: Monitor computing power Monitor computing power of an ML model in the long-term Energy efficiency (Cao et al., 2021)(Kumar et al., 2020)
@@ 4.7. Targeted Quality Attributes
Energy efficiency The ability to accomplish a task while minimizing energy consumption 21
Performance The efficiency with which a task is achieved (e.g., speed, stability) 3
Accuracy The level of how accurately the algorithm performs specified tasks. 3
Recoverability The ability to restore and resume normal operations after a failure 2
Resource utilization The ability to use and allocate resources efficiently 1
@@ Acknowledgements.
We kindly thank our three focus group experts for their valuable time and feedback! Additionally, we thank Iffat Fatima (VU Amsterdam) for her support with integrating the tactics into the AADT. Lewis and Ozkaya’s work was supported by the Department of Defense under Contract No. FA8702-15-D-0002 with Carnegie Mellon University for the operation of the Software Engineering Institute, a federally funded research and development center (DM23-2363). Muccini’s work was supported by the PNRR ICSC National Research Centre for High Performance Computing, Big Data and Quantum Computing (CN00000013), under the NRRP MUR program funded by the NextGenerationEU. This research was supported by ExtremeXP, a project co-funded by the European Union Horizon Programme under Grant Agreement No. 101093164.